# ejcolrgb1,0,0 The MandelZoom project II: the impact of stellar feedback on black hole accretion through an -disc in dwarf galaxies with a resolved interstellar medium

### v1

@@ 3.4.2 Impact of early stellar feedback on the CND: increased susceptibility to SNe
As shown in the first columns of the respective runs, which is 8 Myr before the onset of SNe, the stellar distribution in both simulations lies mainly near the CND edge, with only minor differences in the number of stars. However, the gas density and temperature distributions differ markedly. In the SN run, the CND edge remains relatively dense and cold, with K. In contrast, in the full run, massive stars immediately emit EUV photons, heating the CND. As a result, the CND edge where these stars are located is hotter and more diffuse, while gas is compressed inward to – pc from the centre, thereby enhancing BH accretion. This radiative feedback makes the CND more susceptible to subsequent SNe. Despite the CND being slightly less massive in the SN run, it survives throughout the full simulation time of 450 Myr. By contrast, in the full run, once the first SNe explode, the entire CND is disrupted and dispersed, leaving no disc until the next inflow into the nuclear region occurs. Thus, with early stellar feedback, the CND undergoes episodic cycles of destruction and reformation.
@@ 4 Discussion
The studies most closely related to ours are those of Partmann et al. (2025) and Petersson et al. (2025). Our results are in broad agreement with Partmann et al. (2025), who model dwarf galaxies with an NSC, but without additional refinement around the BH, where, similarly to our work, they find that the presence of an NSC promotes gas accretion onto IMBHs (see Paper I for a more detailed comparison). Petersson et al. (2025) modelled IMBH accretion in isolated dwarf galaxies with a subgrid accretion disc model, but without incorporating an NSC. In their simulations, BH accretion occurred intermittently even in models including only SN feedback, whereas in our simulations, the CND was able to withstand SN feedback and survive, leading to continuous BH accretion over 450 Myr. We attribute these different outcomes to the absence of an NSC in their models and the adoption of a different stellar feedback injection scheme, which by construction preferentially couples feedback into the dense surrounding gas, rather than isotropically (see e.g., Hopkins et al., 2018; Smith et al., 2018a). Hence, in Petersson et al. (2025), the gas near the BH may not be sufficiently dense and is thus easily disrupted by SN explosions, while in our case, despite SN events occurring tens of times more frequently within the central pc region, the CND remained structurally intact.
@@
In the remaining paragraphs, we highlight several limitations of our model and outline possible directions for future improvements (see also Paper I). The most important limitation is that, for the sake of simplicity, we focused exclusively on the impact of stellar feedback on BH accretion, while AGN feedback was not included. However, AGN feedback, such as jet, wind and radiation feedback, can strongly affect the CND and the ISM, and consequently regulate the inflow rate of gas into the galactic centre (see e.g., Bourne et al., 2014, 2015; Mukherjee et al., 2018; Torrey et al., 2020; Talbot et al., 2022; Mercedes-Feliz et al., 2023). It remains to be understood if AGN feedback is able to significantly affect the local ISM and CND properties, and hence the mass supply onto the central BH, and in our future work, we will tackle this crucial issue.
Our accretion disc model follows a radiatively efficient -disc prescription, appropriate for geometrically thin and optically thick flows. However, when the CND is disrupted or dispersed and the accretion rate drops (e.g. with ), the flow may transition toward a radiatively inefficient, geometrically thick state (e.g., Yuan & Narayan, 2014; Koudmani et al., 2024). This transition not only affects the BH mass growth, but also the spin magnitude and direction evolution, which is especially important to model accurately once AGN jet feedback is considered. A more detailed treatment of this physical regime will be investigated in future extensions of this study.
@@
Recent studies, particularly in the IMBH mass regime, suggest that BHs embedded in NSCs can grow efficiently through the TDE channel (see e.g., Zubovas, 2019; Alexander & Bar-Or, 2017; Rizzuto et al., 2023; Chang et al., 2025; Rantala et al., 2025). In our model, however, the NSC was represented with a single-mass population experiencing softened gravitational interactions ( pc), which prevents us from capturing the TDE channel as well as processes such as relaxation and mass segregation within the NSC. It will be particularly timely to consider IMBH growth through TDEs in full galaxy formation simulations, which, if significant, will alleviate the difficulty of feeding (low mass) IMBH within shallow or gas-poor gravitational potentials.
@@ Acknowledgements
We would like to thank Antti Rantala and Thorsten Naab for their valuable discussions and insightful feedback on this manuscript. E.S. and D.S. acknowledge support from the Science and Technology Facilities Council (STFC) under grant ST/W000997/1. M.A.B. is supported by a UKRI Stephen Hawking Fellowship (EP/X04257X/1). S.K. acknowledges support from an 1851 Research Fellowship awarded by the Royal Commission for the Exhibition of 1851, as well as a Junior Research Fellowship at St Catharine’s College, Cambridge. This work was enabled by the University of Cambridge Research Computing Service (www.csd3.cam.ac.uk), funded by Dell EMC and Intel using Tier2 funding from the Engineering and Physical Sciences Research Council (capital grant EP/P020259/1), together with DiRAC funding from the STFC (www.dirac.ac.uk). We also made use of the DiRAC@Durham facility managed by the Institute for Computational Cosmology on behalf of the STFC DiRAC HPC Facility. The equipment was funded by BEIS capital funding via STFC capital grants ST/P002293/1, ST/R002371/1 and ST/S002502/1, Durham University, and STFC operations grant ST/R000832/1. DiRAC forms part of the National e-Infrastructure.

### v2

@@ 3.4.2 Impact of early stellar feedback on the CND: increased susceptibility to SNe
[…]
@@ 4 Discussion
The studies most closely related to ours are those of Partmann et al. (2025) and Petersson et al. (2025). Our results are in broad agreement with Partmann et al. (2025), who model dwarf galaxies with an NSC, but without additional refinement around the BH, where, similarly to our work, they find that the presence of an NSC promotes gas accretion onto IMBHs (see Paper I for a more detailed comparison). Petersson et al. (2025) modelled IMBH accretion in isolated dwarf galaxies with a subgrid accretion disc model, but without incorporating an NSC. In their simulations, BH accretion occurred intermittently even in models including only SN feedback, whereas in our simulations, the CND was able to withstand SN feedback and survive, leading to continuous BH accretion over 450 Myr. We attribute these different outcomes to the absence of an NSC in their models and the adoption of a different stellar feedback injection scheme, which by construction preferentially couples feedback into the dense surrounding gas, rather than isotropically (see e.g., Hopkins et al., 2018; Smith et al., 2018b). Hence, in Petersson et al. (2025), the gas near the BH may not be sufficiently dense and is thus easily disrupted by SN explosions, while in our case, despite SN events occurring tens of times more frequently within the central pc region, the CND remained structurally intact.
@@
ejcolOver the course of the 450 Myr simulation, the NSC mass grows only marginally through intermittent in-situ star formation, by (approximately –), while the BH mass increases by – (–). This differential growth leads to a modest upward shift in the plane, yet the evolutionary track remains well within the intrinsic scatter of the observed relation (e.g., Nguyen et al., 2018). In terms of NSC structural evolution, the effective radius of the NSC shows no significant oscillations or systematic trends in most runs over 450 Myr.
[…]
ejcolHowever, AGN feedback, such as jet, wind and radiation feedback, can strongly affect the CND and the ISM, and consequently regulate the inflow rate of gas into the galactic centre (see e.g., Bourne et al., 2014, 2015; Mukherjee et al., 2018; Torrey et al., 2020; Talbot et al., 2022; Mercedes-Feliz et al., 2023). It remains to be understood if AGN feedback is able to significantly affect the local ISM and CND properties, and hence the mass supply onto the central BH, and in our future work, we will tackle this crucial issue.
Our accretion disc model follows a radiatively efficient -disc prescription, appropriate for geometrically thin and optically thick flows. However, when the CND is disrupted or dispersed and the accretion rate drops (e.g. with ), the flow may transition toward a radiatively inefficient, geometrically thick state (e.g., Yuan and Narayan, 2014; Koudmani et al., 2024). This transition not only affects the BH mass growth, but also the spin magnitude and direction evolution, which is especially important to model accurately once AGN jet feedback is considered. A more detailed treatment of this physical regime will be investigated in future extensions of this study.
@@
Recent studies, particularly in the IMBH mass regime, suggest that BHs embedded in NSCs can grow efficiently through the TDE channel (see e.g., Zubovas, 2019; Alexander and Bar-Or, 2017; Rizzuto et al., 2023; Chang et al., 2025; Rantala et al., 2025). In our model, however, the NSC was represented with a single-mass population experiencing softened gravitational interactions ( pc), which prevents us from capturing the TDE channel as well as processes such as relaxation and mass segregation within the NSC. It will be particularly timely to consider IMBH growth through TDEs in full galaxy formation simulations, which, if significant, will alleviate the difficulty of feeding (low mass) IMBH within shallow or gas-poor gravitational potentials.
@@ Acknowledgements
ejcolWe thank an anonymous referee for the valuable suggestions that have helped to improve our paper. We also would like to thank Antti Rantala and Thorsten Naab for their valuable discussions and insightful feedback on this manuscript. E.S. and D.S. acknowledge support from the Science and Technology Facilities Council (STFC) under grant ST/W000997/1. M.A.B. is supported by a UKRI Stephen Hawking Fellowship (EP/X04257X/1). S.K. acknowledges support from an 1851 Research Fellowship awarded by the Royal Commission for the Exhibition of 1851, as well as a Junior Research Fellowship at St Catharine’s College, Cambridge. This work was enabled by the University of Cambridge Research Computing Service (www.csd3.cam.ac.uk), funded by Dell EMC and Intel using Tier2 funding from the Engineering and Physical Sciences Research Council (capital grant EP/P020259/1), together with DiRAC funding from the STFC (www.dirac.ac.uk). We also made use of the DiRAC@Durham facility managed by the Institute for Computational Cosmology on behalf of the STFC DiRAC HPC Facility. The equipment was funded by BEIS capital funding via STFC capital grants ST/P002293/1, ST/R002371/1 and ST/S002502/1, Durham University, and STFC operations grant ST/R000832/1. DiRAC forms part of the National e-Infrastructure.